# POPGym Arcade: Parallel Pixelated POMDPs

### v1

@@ 5 Experiments
We seek to answer the following questions through experiments. See Appendices A, C, B and D for additional results and Appendices G and H for architecture and hyperparameter details. All plot shading represents the 95% bootstrapped confidence interval.
@@
On one hand, we leverage compilation and vectorization, but on the other hand, we have large pixel observation spaces. We compare the throughput of our environments to other well-known environments over ten random seeds in Fig. 2 on an RTX4090 GPU and Intel I7 13700 CPU.
@@
Figure 2: We plot the throughput of our environments in frames per second (FPS), and compare to prior work. We achieve roughly linear scaling with the number of environments, but eventually saturate GPU cores, resulting in sublinear scaling. Still, our method scales better than Atari, and outperforms MinAtar and POPGym benchmarks while producing observations that are orders of magnitude larger.
How does the MinGRU fare in online value estimation?
Prior work finds that the MinGRU routinely outperforms SSMs in a decision transformer on offline RL tasks. Does this finding generalize to online RL? In Fig. 3(a), we compare the LRU to the MinGRU across all tasks, and provide a breakdown by observability.
Model MDP POMDP Total
MLP $\mathbf{0.57}\pm 0.02$ $0.33\pm 0.01$ $0.45\pm 0.01$
MinGRU $0.38\pm 0.02$ $0.45\pm 0.03$ $0.42\pm 0.02$
LRU $0.48\pm 0.06$ $\mathbf{0.59}\pm 0.01$ $\mathbf{0.53}\pm 0.04$
(a) The return over all tasks, variants, and difficulties – grouped by model and observability, and reduced using the mean. We represent the standard deviation between seeds. Surprisingly, the MinGRU and LRU tend to perform better on POMDPs than on MDPs.
(b) For BattleShip, both MinGRU and LRU provide better sample efficiency under partial observability than full observability.
Figure 3: Comparing POMDP-MDP twins, see Fig. 1 for a visual comparison of environment twins.
(a) (MDP) When presented with Markov observations, the policy tends to rely on prior observations to estimate the Q value. The policy bases the Q value on all prior cursor locations, and relies more heavily on the “hit” marker in $o_{2},o_{4}$ than it does at $o_{9}$, the current timestep.
(b) (POMDP) The POMDP-trained policy behaves similarly to the MDP, recalling prior hits and misses to estimate the Q value. Here, pixel activations are fewer than the MDP.
Figure 4: Visualization of the LRU, produced by our memory analysis tools. Memory-endowed policies tend to ignore the Markov property.
What role does memory play? Does observability impact behavior?
We investigate how our policies utilize memory. With pixel-space observations, we can draw conclusions about how pixels persist in memory, and their impact on Q value predictions. For this experiment, we train for 100M environment steps, until policies converge (Fig. 13). Then, we use the gradient analysis tools from Section 4.3 to inspect how individual memories impact Q values. We plot the partial derivatives for BattleShip in Fig. 4 (see Appendix B for more tasks).
6 Discussion
@@
Our environments are faster than CPU-based environments and on-par with other GPU-based environments that have much smaller observation spaces. For example, our observations are five orders of magnitude larger than POPGym, yet we run approximately one order of magnitude faster.
SSMs outperform the MinGRU in online value estimation.
Although Feng et al. (2024) find that the MinGRU outperforms SSMs in the offline decision transformer setting, our experiments in online value estimation tasks reveal that the LRU SSM delivers superior performance (Fig. 3(a)). The MLP outperforms both memory models on MDPs, suggesting that memory-endowed policies find suboptimal solutions in the exponentially larger POMDP policy space.
Training memory models on POMDPs can be more efficient than MDPs.
In theory, solving POMDPs requires exponentially more training samples than MDPs (Kaelbling et al., 1998), but our empirical evidence suggests otherwise. On average, both the LRU and MinGRU demonstrate slightly better sample efficiency on POMDPs compared to MDPs (Fig. 3(a)). We drill down into specific tasks in Fig. 3(b). We initially suspected POMDP observation aliasing increased policy churn thereby improving exploration (Schaul et al., 2022), but we found this was not the case (Fig. 14). We provide further explanation in the following paragraph.
Memory models ignore the Markov property.
Rather than consider POMDPs easy, we should instead consider MDPs difficult. Even in fully observable tasks, recurrent models continue to rely on prior observations to predict Q-values, even when such observations do not affect the return (Figs. 4 and B). For example, Fig. 5 focuses on past agent positions which do not change the return. This leads us to a fairly surprising finding – given memory, observability is a poor predictor of policy performance. Making a task more or less observable often has less effect on the return than difficulty of the fully observable task (Appendix A). SSMs do not have a prior to emphasize more recent observations. As such, Atari and other weak POMDPs could be sufficiently difficult to evaluate such models.
Limitations
We note three limitations of our work. 1. POPGym Arcade rendering logic requires significant GPU memory. We saturate GPU cores with enough parallel environments, resulting in further environments exhibiting sublinear scaling. More powerful GPUs should push this saturation point out further. 2. Fig. 3(a) uses three random seeds. However, deviation is low between seeds, and most analysis occurs over all environments, providing at least 63 datapoints for each cell in Fig. 3(a). 3. We demonstrate that memory models fails to understand the Markov property, but were unable to figure out why this is the case.
@@ 7 Conclusion
We proposed the POPGym Arcade benchmark, a collection of parallel pixel-based POMDPs, MDPs, introspection tools, and associative recurrent memory baselines. Using our library, we tested our environments and found that observability might not be as critical to sample efficiency as previously thought.
Supplementary Materials
The following content was not necessarily subject to peer review.
Appendix A Return Breakdown by Model
Table 2: Return for the LRU model, broken down by task and observability. We also list the absolute difference in return between each MDP/POMDP variant.
Environment MDP Return POMDP Return Absolute Difference
AutoEncodeEasy $0.26\pm 0.00$ $0.26\pm 0.01$ $0.00$
AutoEncodeMedium $\mathbf{0.25}\pm 0.00$ $0.25\pm 0.00$ $0.00$
AutoEncodeHard $\mathbf{0.24}\pm 0.00$ $0.24\pm 0.00$ $0.00$
BattleShipEasy $0.79\pm 0.01$ $\mathbf{0.84}\pm 0.01$ $0.05$
BattleShipMedium $0.44\pm 0.18$ $\mathbf{0.77}\pm 0.02$ $0.33$
BattleShipHard $0.37\pm 0.21$ $\mathbf{0.64}\pm 0.01$ $0.27$
CartPoleEasy $\mathbf{0.99}\pm 0.00$ $0.99\pm 0.00$ $0.00$
CartPoleMedium $\mathbf{0.97}\pm 0.02$ $0.96\pm 0.02$ $0.01$
CartPoleHard $\mathbf{0.96}\pm 0.01$ $0.91\pm 0.02$ $0.05$
CountRecallEasy $\mathbf{0.52}\pm 0.01$ $0.37\pm 0.00$ $0.15$
CountRecallMedium $\mathbf{0.24}\pm 0.03$ $0.24\pm 0.02$ $0.00$
CountRecallHard $0.19\pm 0.01$ $\mathbf{0.21}\pm 0.01$ $0.02$
MineSweeperEasy $\mathbf{0.78}\pm 0.00$ $0.65\pm 0.04$ $0.13$
MineSweeperMedium $\mathbf{0.35}\pm 0.02$ $0.33\pm 0.02$ $0.02$
MineSweeperHard $0.18\pm 0.01$ $\mathbf{0.23}\pm 0.00$ $0.05$
NavigatorEasy $0.10\pm 0.69$ $\mathbf{0.59}\pm 0.01$ $0.49$
NavigatorMedium $-0.36\pm 0.04$ $\mathbf{0.49}\pm 0.02$ $0.85$
NavigatorHard $-0.17\pm 0.07$ $\mathbf{0.48}\pm 0.01$ $0.65$
NoisyCartPoleEasy $\mathbf{0.99}\pm 0.00$ $0.99\pm 0.00$ $0.00$
NoisyCartPoleMedium $\mathbf{0.98}\pm 0.00$ $0.98\pm 0.00$ $0.00$
NoisyCartPoleHard $0.96\pm 0.00$ $\mathbf{0.97}\pm 0.00$ $0.01$
Table 3: Return for the MinGRU model, broken down by task and observability. We also list the absolute difference in return between each MDP/POMDP variant.
Environment MDP Return POMDP Return Absolute Difference
AutoEncodeEasy $0.27\pm 0.00$ $\mathbf{0.28}\pm 0.00$ $0.01$
AutoEncodeMedium $\mathbf{0.26}\pm 0.00$ $0.26\pm 0.00$ $0.00$
AutoEncodeHard $\mathbf{0.25}\pm 0.00$ $0.24\pm 0.01$ $0.01$
BattleShipEasy $0.60\pm 0.07$ $\mathbf{0.68}\pm 0.03$ $0.08$
BattleShipMedium $0.39\pm 0.02$ $\mathbf{0.54}\pm 0.01$ $0.15$
BattleShipHard $0.31\pm 0.03$ $\mathbf{0.43}\pm 0.09$ $0.12$
CartPoleEasy $\mathbf{0.96}\pm 0.00$ $0.89\pm 0.02$ $0.07$
CartPoleMedium $\mathbf{0.90}\pm 0.01$ $0.72\pm 0.02$ $0.18$
CartPoleHard $\mathbf{0.87}\pm 0.09$ $0.53\pm 0.02$ $0.26$
CountRecallEasy $\mathbf{0.40}\pm 0.01$ $0.23\pm 0.01$ $0.17$
CountRecallMedium $\mathbf{0.20}\pm 0.01$ $0.14\pm 0.04$ $0.06$
CountRecallHard $\mathbf{0.17}\pm 0.01$ $0.09\pm 0.03$ $0.08$
MineSweeperEasy $\mathbf{0.71}\pm 0.01$ $0.55\pm 0.02$ $0.16$
MineSweeperMedium $\mathbf{0.29}\pm 0.01$ $0.24\pm 0.00$ $0.05$
MineSweeperHard $0.14\pm 0.01$ $\mathbf{0.15}\pm 0.01$ $0.01$
NavigatorEasy $-0.49\pm 0.02$ $\mathbf{0.48}\pm 0.02$ $0.97$
NavigatorMedium $-0.47\pm 0.02$ $\mathbf{0.36}\pm 0.03$ $0.83$
NavigatorHard $-0.41\pm 0.02$ $\mathbf{0.17}\pm 0.19$ $0.58$
NoisyCartPoleEasy $\mathbf{0.96}\pm 0.01$ $0.90\pm 0.02$ $0.06$
NoisyCartPoleMedium $\mathbf{0.92}\pm 0.02$ $0.85\pm 0.01$ $0.07$
NoisyCartPoleHard $\mathbf{0.85}\pm 0.03$ $0.76\pm 0.01$ $0.09$
@@ Appendix B Additional Memory Analysis
Figure 5: (MDP, Top) The MDP-trained policy focuses on the position of the goal and agent position in prior observations, even though this information is unecessary at $t=9$. In particular, we see the the agent remember explosive tiles from near-death experiences ($o_{2},o_{6}$), where the shortest path would take the agent through explosive tiles. At the risk of anthropomorphizing, one might consider this the emergence of “fear”, or how a traumatic event from the past can persist in memory influence future decision making. (POMDP, Bottom) The agent takes a safer path to the goal. At $o_{9}$, the agent must head south to reach the goal. In $o_{0}$, we see the path from the location at $t=9$ to the goal highlighted, suggesting the agent is planning an unobstructed path to the goal, using information stored at $t=0$.
Figure 6: (MDP, Top) Like in previous experiments, the MDP policy tends to base Q values on the memory corresponding to prior observations of the tiles. We see larger activations when tiles are first revealed at $o_{2},o_{4},o_{6}$ than the equivalent tiles in $o_{9}.$ (POMDP, Bottom) In the POMDP, the agent memory focuses more heavily on the score at the top of the screen. We see hotspots when tiles are first revealed, demonstrating that the policy remembers the location of previously swept tiles.
Figure 7: Partial derivative grid for Battleship MDP
Figure 8: Partial derivative grid for Battleship POMDP
Figure 9: Partial derivative grid for Navigator MDP
Figure 10: Partial derivative grid for Navigator POMDP
Appendix C Return Plots
We present our results as line plots, demonstrating the maximum cumulative episodic reward achieved during each run, separated by environment and the difficulty level (Figs. 11 and 12)
With all the experiments, a sample of 10M timesteps are taken from the respective environment. However, in the case of BattleShip and CountRecall experiments, the sample is set to 20M timesteps.
We use 3 random seed and smooth the line plots using EWMA. The solid area depict the mean and the shaded area is the 95% bootstrap confidence interval.
Figure 11: Return plots (continued)
Figure 12: Return plots (continues on next page)
Figure 13: For our gradient analysis, we train the LRU policies for 100M environment steps until convergence.
Appendix D Policy Churn
Figure 14: We execute a pilot study to see if policy churn is significantly different between policies trained on POMDPs and MDPs. We plot the difference in churn between the MDP and corresponding POMDP policies. For example, if the POMDP policy has a churn ratio of 0.5 (half of states) and the MDP policy has a churn of 0.4, the difference in churn rates would be 0.1. We see negligible differences in churn between POMDPs and MDPs.

### v5

@@ 5 Experiments
Using our benchmark and visualization tools, we seek to answer three questions. See Appendices G and I for architecture and hyperparameter details.
@@
We leverage compilation and vectorization for improved throughput, but our large pixel observations take significant resources to render. In Fig. 2, we compare the throughput of our environments with $\Omega_{\mathrm{small}}$ to other well-known environments over ten random seeds in Fig. 2, using an RTX4090 GPU and Intel I7 13700 CPU. We parallelize CPU environments using synchronous VectorEnvs, and shade the 95% bootstrapped confidence interval.
@@
Figure 2: We plot the throughput of our environments in Frames Per Second (FPS), and compare to prior work. We achieve linear scaling with the number of environments, but saturate GPU cores and decay to sublinear scaling after $2^{7}$ environments. Still, our method scales nearly as well as Gymnax, which uses small non-pixel observation spaces. POPGym Arcade is faster than Atari, and outperforms POPGym and MinAtar benchmarks while producing observations that are orders of magnitude larger.
Figure 3: We plot the normalized returns $\in[0,1]$ over 5 seeds and all observability, memory model, and difficulty configurations for all environments. The boxes, line, and dot represent the quartiles, median, and mean respectively. We organize the results by observability and model for a counterfactual study of memory.
(a) Memory saliency map for the MineSweeper MDP.
(b) Memory saliency map for the MineSweeper POMDP.
Figure 4: How do our fully-trained agents use memory? We plot $o_{0},\dots,o_{9}$ in the upper row. In the lower row, we plot pixel-wise recall (Eq. 4) to compute $Q(\hat{s}_{9},a_{9})$ with respect to each observation. Surprisingly, LRU policies ignore the Markov property, relying on older observations instead. Note that our policy architecture utilizes a skip connection across the recurrent model, allowing the policy to bypass memory if not necessary.
To what extent does memory mitigate partial observability?
We execute a large scale study, training policies on all proposed environment, difficulty, and observability configurations over five random seeds. We evaluate three different associative recurrent backbones and an MLP backbone, and train the policy and backbones end-to-end using PQN. Using MDP performance as the control, we can measure exactly how well each memory architecture mitigates partial observability. We provide the full results in Appendix D, and aggregate results by model and observability in Fig. 3.
What role does memory play in RL?
We investigate how our trained policies utilize memory. With pixel-space observations, we can visualize which pixels our agent recalls to make decisions. First, we use the memory analysis tools from Section 4.2 to inspect how individual memories qualitatively impact $Q$ values (Figs. 4 and B). Then, we compute recall densities (Definition 4.1) over environment and model configurations to produce Appendix A. We aggregate densities across models into an environment summary in Fig. 5. Finally, we execute a toy experiment that introduces out-of-distribution observations into a recurrent policy, and investigate the results (Figs. 6 and C). In particular, we inject Gaussian noise $\mathcal{N}(0,1)$ into an observation, and determine how this affects future $Q$ values and actions for the LRU policy. Note that we generate a fixed, noise-free trajectory first using the trained policy. Only after do we inject noise into the trajectory, ensuring that the noise alone is responsible for changes in $Q$ value.
Figure 5: A quantitative inspection of memory. We use Definition 4.1 to estimate LRU recall density $\mathbb{E}_{\pi,f}[\delta_{Q}(\mathbf{x},\tau)]$ for binned $\tau$ representing the start, middle, and end of a trajectory. We estimate the density using five trajectories per memory model, and aggregate over all memory models. The recall density tells us which observations the agent recalls for decision making. If memory models learned the Markov property (to ignore older observations), we would expect all MDP density to be in the $1.00.66\leq\tau<$ 1.0 bin. This is not what we see – in all cases, the distribution is spread throughout the full trajectory. Even for simple tasks like CartPole, the agent places considerable density on useless observations. Dependence on many prior observations suggests the policies may not generalize well.
Figure 6: We can poison the recurrent state with out-of-distribution observations. First, we generate a trajectory following the greedy policy. Then, we inject noise $\epsilon$ into one selected observation, and examine how this information propagates into future actions. Note that each row follows an identical trajectory up to $t=10$, and noise is only present in one observation for each row. We see that not only does the future $Q$ value change significantly, but so does the greedy action.
6 Results
@@
Our environments are faster than all CPU-based environments we tested (Fig. 2). For example, our observations are four orders of magnitude larger than POPGym, yet we run approximately two orders of magnitude faster. We achieve similar throughput to hardware-accelerated Gymnax while generating observations four orders of magnitude larger. One can quickly prototype models or algorithms on our pixel-space, partially observable tasks.
Memory models introduce confounding factors.
There is a noticeable performance gap between memory models in our experiments (Fig. 3). The large gap between the POMDP and MDP control for the MLP demonstrates that memory is necessary to solve our benchmark.
Without confounding factors, all memory models should have the same MDP performance as the MLP, which is not the case. We initially suspected that the parameter increase associated with recurrent models would be a positive confounding factor, artificially boosting the final return compared to the MLP. The sizable gaps (MLP/MDP - FART/MDP) and (MLP/MDP - MinGRU/MDP) instead suggest large negative confounding factors – adding FART or MinGRU to a policy that does not need it is more likely to harm performance rather than help. Unlike the FART or MinGRU, the LRU appears to have marginal positive confounding factors (MLP/MDP - LRU/MDP). These findings cast doubt on prior work without control MDPs, emphasizing the importance of environment twins.
Recurrent policies ignore the Markov property.
Figs. 4, B, 5 and A all tell the same story – recurrent models tend to solve POMDPs and MDPs in the same manner, ignoring the Markov property. Even in fully observable tasks, recurrent models continue to rely on prior observations to predict $Q$ values, even when such observations do not affect the return. We stress that all our recurrent policies include a residual connection that bypasses the memory model, allowing the agent to ignore prior observations if not necessary.
Out-of-distribution observations poison recurrent states.
Recurrent policies rely on many unnecessary prior observations (Figs. 4 and 5). Therefore, out-of-distribution observations can “poison” the recurrent state, affecting $Q$ values and agent decisions far into the future (Figs. 6 and 12). Both $Q$ values and policy actions are sensitive to this phenomena. Demonstrating this on a simple task like CartPole does not bode well for more complex and realistic tasks. These results have far-reaching implications for offline RL, imitation learning, or sim-to-real scenarios, where we expect the agent to encounter out-of-distribution observations after training. A single out-of-distribution observation could poison the agent, altering its behavior long after the observation occurred.
6.1 Limitations and Future Work
Should we limit context?
Recurrent models exhibit poor robustness due to reliance on many past observations. Unlike recurrent policies, a framestack or transformer policy may be harder to poison. They will discard the out-of-distribution observation when it exceeds the context window, returning to an in-distribution context.
Regularization may push parameters towards to simpler policies that rely on fewer observations, but one must be careful combining standard regularization with recurrent models. For example, applying naive $L_{2}$ regularization could worsen the problem for the MinGRU, by minimizing $W_{1},b_{1}$ in the forget gate and retaining observations for longer: $h_{t}=(1-\sigma(W_{1}x_{t}+b_{1}))\odot h_{t}+\dots$. Recurrent dropout or similar methods may be better suited to the task.
Finally, our poisoning experiment is just a proof of concept, not a rigorous analysis. Non-toy problems have more complex observation models that could impact $Q$ values differently.
Resource constraints
POPGym Arcade rendering logic eventually saturates GPU cores given enough parallel environments, resulting in further environments exhibiting sublinear scaling (Fig. 2). More powerful GPUs should push this saturation point out further.
We have insufficient resources to perform proper grid searches for model-specific hyperparameters, given the number of environment configurations. Instead, we slightly modify the parameters from the PQN paper to reduce the return variance between random seeds. Further tuning could improve the performance of various recurrent models.
@@ 7 Conclusion
We proposed POPGym Arcade, a collection of hardware-accelerated, pixel-based POMDPs, MDPs, and memory introspection tools. Using our library, we detected powerful confounding factors when evaluating memory models. We found that memory models use unnecessary information to build Markov states, with negative consequences for robustness.
8 Acknowledgements
We are thankful to the University of Macau for supporting this work through grant SRG2025-00006-FST.
Appendix
Appendix A Recall Density Breakdown by Model and Task
In this section, we plot the recall density after training, for each model and task. We discretize $\tau$ into sixteen bins and compute the density over five random seeds. No matter the memory model, uninformative prior observations still tend to affect future decisions in MDPs. Even in the best case (CartPole, MinGRU), roughly half the density is distributed across old, uninformative observations (Fig. 9).
Figure 7: Recall density of FART categorized by environment, over five random seeds.
Figure 8: Recall density of LRU categorized by environment, over five random seeds.
Figure 9: Recall density of MinGRU categorized by environment, over five random seeds.
@@ Appendix B Additional Memory Analysis
We provide further pixel-level analysis of our environments using Eq. 4. We see that memory models ignore Markov observations, still storing and recalling information in the MDP case.
(a) Memory saliency map for the Battleship MDP
(b) Memory saliency map for the Battleship POMDP
Figure 10: LRU saliency plotted via Eq. 4.
(a) Memory saliency map for the Navigator MDP
(b) Memory saliency map for the Navigator POMDP
Figure 11: LRU saliency plotted via Eq. 4.
Appendix C Recurrent Policy Poisoning
We provide additional experiments for recurrent policy poisoning. We use the cartpole example for even longer rollout horizons of 50 and 100 timesteps. We find that old observations still have a significant impact on the $Q$ values and corresponding actions.
Figure 12: We demonstrate recurrent policy poisoning over longer horizons, using the LRU model. At 50 timesteps in the future, corrupted observations can still influence $Q$ values enough to change the selected action. At 100 timesteps, adding noise results in negative $Q$ values, but does not change the action in this specific case.
Appendix D Return Breakdown by Model
In this section, we provide the unnormalized returns (mean and standard deviation) for the aggregated results in Fig. 3. We provide values for each environment, memory model, and difficulty level.
Table 2: Return for the LRU model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy 0.26 ± 0.00 0.27 ± 0.01
AutoEncodeMedium 0.24 ± 0.01 0.24 ± 0.00
AutoEncodeHard 0.23 ± 0.02 0.22 ± 0.02
BattleShipEasy 0.85 ± 0.02 0.84 ± 0.01
BattleShipMedium 0.78 ± 0.03 0.82 ± 0.01
BattleShipHard 0.73 ± 0.04 0.8 ± 0.01
CartPoleEasy 0.99 ± 0.00 0.99 ± 0.00
CartPoleMedium 0.98 ± 0.01 0.98 ± 0.01
CartPoleHard 0.97 ± 0.01 0.95 ± 0.02
CountRecallEasy 0.47 ± 0.06 0.36 ± 0.01
CountRecallMedium 0.22 ± 0.03 0.23 ± 0.04
CountRecallHard 0.15 ± 0.02 0.23 ± 0.01
MineSweeperEasy 0.87 ± 0.00 0.82 ± 0.01
MineSweeperMedium 0.67 ± 0.01 0.66 ± 0.01
MineSweeperHard 0.62 ± 0.01 0.61 ± 0.00
NavigatorEasy 0.91 ± 0.01 0.44 ± 0.09
NavigatorMedium 0.91 ± 0.01 0.32 ± 0.03
NavigatorHard 0.94 ± 0.01 0.4 ± 0.06
NoisyCartPoleEasy 0.99 ± 0.00 0.99 ± 0.00
NoisyCartPoleMedium 0.98 ± 0.01 0.99 ± 0.00
NoisyCartPoleHard 0.97 ± 0.01 0.97 ± 0.01
Table 3: Return for the MinGRU model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy 0.27 ± 0.00 0.27 ± 0.00
AutoEncodeMedium 0.25 ± 0.01 0.26 ± 0.00
AutoEncodeHard 0.25 ± 0.00 0.25 ± 0.00
BattleShipEasy 0.82 ± 0.02 0.76 ± 0.01
BattleShipMedium 0.73 ± 0.02 0.75 ± 0.02
BattleShipHard 0.71 ± 0.03 0.73 ± 0.02
CartPoleEasy 0.98 ± 0.01 0.87 ± 0.02
CartPoleMedium 0.93 ± 0.02 0.68 ± 0.06
CartPoleHard 0.88 ± 0.08 0.54 ± 0.03
CountRecallEasy 0.31 ± 0.01 0.29 ± 0.03
CountRecallMedium 0.17 ± 0.00 0.21 ± 0.04
CountRecallHard 0.15 ± 0.01 0.13 ± 0.04
MineSweeperEasy 0.83 ± 0.01 0.76 ± 0.01
MineSweeperMedium 0.64 ± 0.00 0.61 ± 0.00
MineSweeperHard 0.58 ± 0.01 0.58 ± 0.01
NavigatorEasy 0.89 ± 0.02 0.33 ± 0.08
NavigatorMedium 0.74 ± 0.08 0.25 ± 0.12
NavigatorHard 0.13 ± 0.45 0.08 ± 0.08
NoisyCartPoleEasy 0.95 ± 0.01 0.89 ± 0.01
NoisyCartPoleMedium 0.92 ± 0.01 0.86 ± 0.01
NoisyCartPoleHard 0.86 ± 0.02 0.81 ± 0.01
Table 4: Return for the FART model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy 0.25 ± 0.00 0.26 ± 0.00
AutoEncodeMedium 0.23 ± 0.01 0.23 ± 0.01
AutoEncodeHard 0.2 ± 0.03 0.21 ± 0.02
BattleShipEasy 0.66 ± 0.09 0.77 ± 0.02
BattleShipMedium 0.51 ± 0.01 0.52 ± 0.02
BattleShipHard 0.5 ± 0.00 0.5 ± 0.00
CartPoleEasy 0.98 ± 0.01 0.97 ± 0.01
CartPoleMedium 0.85 ± 0.06 0.78 ± 0.05
CartPoleHard 0.62 ± 0.06 0.6 ± 0.04
CountRecallEasy 0.39 ± 0.03 0.37 ± 0.02
CountRecallMedium 0.06 ± 0.00 0.1 ± 0.04
CountRecallHard 0.06 ± 0.01 0.05 ± 0.00
MineSweeperEasy 0.76 ± 0.01 0.73 ± 0.04
MineSweeperMedium 0.61 ± 0.02 0.59 ± 0.02
MineSweeperHard 0.55 ± 0.00 0.55 ± 0.01
NavigatorEasy 0.46 ± 0.14 0.36 ± 0.05
NavigatorMedium -0.27 ± 0.19 -0.28 ± 0.19
NavigatorHard -0.45 ± 0.02 -0.41 ± 0.02
NoisyCartPoleEasy 0.98 ± 0.01 0.97 ± 0.02
NoisyCartPoleMedium 0.96 ± 0.02 0.96 ± 0.01
NoisyCartPoleHard 0.92 ± 0.02 0.89 ± 0.01
Table 5: Return for the MLP model, broken down by task and observability.
Environment MDP Return POMDP Return
AutoEncodeEasy 0.26 ± 0.00 0.27 ± 0.00
AutoEncodeMedium 0.25 ± 0.00 0.25 ± 0.00
AutoEncodeHard 0.25 ± 0.00 0.25 ± 0.00
BattleShipEasy 0.85 ± 0.02 0.6 ± 0.04
BattleShipMedium 0.7 ± 0.04 0.56 ± 0.01
BattleShipHard 0.75 ± 0.03 0.53 ± 0.00
CartPoleEasy 0.99 ± 0.00 0.86 ± 0.02
CartPoleMedium 0.96 ± 0.01 0.47 ± 0.01
CartPoleHard 0.95 ± 0.01 0.31 ± 0.01
CountRecallEasy 0.4 ± 0.04 0.15 ± 0.01
CountRecallMedium 0.17 ± 0.02 0.06 ± 0.00
CountRecallHard 0.15 ± 0.01 0.08 ± 0.00
MineSweeperEasy 0.86 ± 0.01 0.72 ± 0.01
MineSweeperMedium 0.65 ± 0.01 0.59 ± 0.00
MineSweeperHard 0.57 ± 0.00 0.55 ± 0.00
NavigatorEasy 0.88 ± 0.02 -0.2 ± 0.05
NavigatorMedium 0.86 ± 0.01 -0.19 ± 0.02
NavigatorHard 0.94 ± 0.01 -0.07 ± 0.08
NoisyCartPoleEasy 0.94 ± 0.01 0.82 ± 0.02
NoisyCartPoleMedium 0.86 ± 0.01 0.74 ± 0.01
NoisyCartPoleHard 0.7 ± 0.01 0.66 ± 0.01